2x^2 + 5x + 5 = 0
I have to use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.

Answer by stanbon(75887)

2x^2 + 5x + 5 = 0
b^2-4ac=25-4(2)(5)<0
Therefore there are no Real Number roots.
There are two Complex Number roots.
Cheers,
Stan H.